# Learning Massive-scale Partial Correlation Networks in Clinical Multi-omics Studies with HP-ACCORD

### v1

@@ Abstract
Ultrahigh-dimensional molecular data,
Multi-modal data,
High-performance statistical computing,
Graphical model selection,
Pseudolikelihood,
Communication-avoiding linear algebra,
keywords:
[…]
, and
@@ 1 Introduction
Omics technologies are essential discovery tools for the analysis of complex biological samples such as tissues and biofluids. Individual omics modalities can generate high-dimensional data (large $p$) for characterizing complex biological samples in a high-throughput manner (large $n\ll p$), and recently for deconvoluting molecular profiles of isolated single cells or spatially connected micrometer-sized regions over tissue sections (Vandereyken et al., 2023). Given the unprecedented opportunity to assemble multi-omic data as fine-resolution descriptors of gene expression regulation, cellular metabolism and signal transduction in a biological system, clinical and basic research studies routinely employ a combination of omics platforms to characterize joint variation in the genome, epigenome, transcriptome, proteome and metabolome (Hasin, Seldin and Lusis, 2017; Huang, Chaudhary and Garmire, 2017).
@@
In omics-scale analysis, biological network is typically represented as an adjacency matrix of data features based on marginal pairwise correlations. In the popular weighted gene co-expression network analysis (WGCNA, Zhang and Horvath, 2005), for example, Pearson correlation-based network modules are captured first and the initial network is refined by topological analysis. Despite the success of WGCNA and other extensions in identifying gene communities from omics data, the adjacency of two molecular features as determined by marginal correlations can be a byproduct of shared regulatory factors rather than direct biological interaction. Hence it is natural to estimate the co-expression network underlying a gene expression dataset with a graph representing the conditional dependencies, e.g., partial correlations. The same logic extends to the inference of other types of biological networks from multi-omic data such as metabolic pathways and signaling cascades.
@@
It is well known that the precision matrix, or inverse covariance matrix, of a multivariate distribution encodes partial correlations. The most popular high-dimensional precision matrix estimation approach is the graphical lasso (Yuan and Lin, 2007; Friedman, Hastie and Tibshirani, 2008), which maximizes a multivariate Gaussian likelihood with an $\ell_{1}$-penalty. However, at the scale of modern multi-omics datasets, it is computationally infeasible to estimate desired networks using this strategy mainly due to the computational bottleneck, which is not easily resolved by using a more powerful computer. As we demonstrate later, our application data features close to 300,000 variables. On two high-performance computing (HPC) environments, the most scalable implementation of graphical lasso, BigQUIC (Hsieh et al., 2013), could complete the computation only for the trivial cases in which the computed precision matrix estimates were diagonal. Similarly, a fast implementation (Pang, Liu and Vanderbei, 2014) of CLIME (Cai, Liu and Luo, 2011), another popular methodology, was also not applicable on a HPC machine with 192GB of memory when the dataset dimension exceeded 30,000. In the next section, we discuss in detail that the inherent nature of the optimization problems from these popular methods makes it difficult to scale up to modern omics-scale problems.
@@
Feature screening procedures, which reduces the number of variables before the network estimation step, have been proposed as a remedy to address the lack of scalability of methods mentioned above (Luo, Song and Witten, 2014; Zheng et al., 2020). Although feature screening could be suitable when a moderate dimensionality reduction is necessary, an order of magnitude reduction (300,000 or more to 30,000 or fewer features) precludes functionally important molecules from being considered for network estimation for computational tractability rather than a biologically motivated rationale. It is also difficult to discern essential variables and dispensable ones a priori solely based on numerical criteria chosen by an analyst, hence there is a clear merit in estimating the model with all possible input variables included from the start.
@@
Another category of graphical model structure learning approach is based on optimizing an $\ell_{1}$-penalized pseudolikelihood-based objective function (Peng et al., 2009; Khare, Oh and Rajaratnam, 2015). Although the structure of pseudolikelihood-based objectives can be advantageous from an optimization perspective (Koanantakool et al., 2018), existing estimators do not possess the statistical properties desirable for multi-omics analyses, as we also discuss in the next section.
@@ 2 Challenges
where $\partial\|\boldsymbol{\Theta}\|_{1}$ denotes the subdifferential of the convex function $\boldsymbol{X}\mapsto\|\boldsymbol{X}\|_{1}$ at $\boldsymbol{\Theta}$. The vast literature on the algorithms for graphical lasso, e.g., d’Aspremont, Banerjee and El Ghaoui (2008); Friedman, Hastie and Tibshirani (2008); Li et al. (2010); Hsieh et al. (2011, 2014), essentially reduces to how to solve the KKT equation (2.1) iteratively.
@@
Consequently, most existing algorithms resort to employing an (inner) iterative method to compute $\boldsymbol{\Theta}^{-1}$. For instance, Friedman, Hastie and Tibshirani (2008) invoke a $p$-dimensional lasso regression solver $p$ times every (outer) iteration to compute one column of $\boldsymbol{\Theta}^{-1}$ at a time, which becomes prohibitively expensive in omics-scale problems. BigQUIC (Hsieh et al., 2013) updates $\boldsymbol{\Theta}$ by block coordinate descent and then computes $\boldsymbol{\Theta}^{-1}$ column by column, by storing only a small fraction of the $p$ columns of $\boldsymbol{\Theta}^{-1}$ in a cache and recomputing the missing columns on demand; the recomputation solves a fraction of the $p$ linear equations defining the inverse ($\boldsymbol{\Theta}\boldsymbol{X}=\boldsymbol{I}_{p}$) by using conjugate gradient. Since this strategy is only successful if the cache miss rate is low, BigQUIC partitions $\boldsymbol{\Theta}$ into a block matrix and permutes blocks to minimize the number of the off-block-diagonal elements. Therefore, the scalability of BigQUIC is limited by the maximum degree of the underlying graph. For more complex graph structures of interest, e.g., those that arise in multi-omics studies, computable dimensions are practically less than $100,000$; see §5.
CLIME minimizes $\|\boldsymbol{\Theta}\|_{1}$ subject to the constraint $\|\boldsymbol{S}\boldsymbol{\Theta}-\boldsymbol{I}_{p}\|_{\infty}\leq\lambda$ over $\mathbb{R}^{p\times p}$, where $\|\cdot\|_{\infty}$ is the vector $\ell_{\infty}$ norm. The optimization problem of CLIME can be decomposed into $p$ independent, $p$-dimensional linear programming (LP) problems, each solving a column of $\boldsymbol{\Theta}$. While LP is the most extensively studied convex optimization problem, commercial LP solvers like Gurobi or Cplex, which utilize interior-point or simplex methods, struggle with scalability when handling hundreds of thousands of variables (Mazumder, Wright and Zheng, 2019). Exacerbating the situation, for omics-scale CLIME we are required to solve hundreds of thousands of large-scale LP problems. This remains an almost unattainable task despite the independence. As a result, fastclime (Pang, Liu and Vanderbei, 2014; Pang et al., 2017), which could not handle more than 30,000 variables in our problem instance, remains one of the most scalable implementations publicly available.
As the most scalable pseudolikelihood-based method up to date, CONCORD (Khare, Oh and Rajaratnam, 2015) solves the $\ell_{1}$-penalized minimization problem
@@
where $\boldsymbol{\Theta}_{D}$ denotes the diagonal and $\boldsymbol{\Theta}_{X}$ denotes the off-diagonal parts of $\boldsymbol{\Theta}$. It is a convex amendment of a non-convex pseudolikelihood-based method SPACE (Peng et al., 2009), which in turn extends the node-wise regression approach by Meinshausen and Bühlmann (2006) to estimate $\boldsymbol{\Theta}^{*}$. [^1] CLIME can be understood as an extension of the latter in a different direction, with the aforementioned scalability bottleneck. The associated KKT condition is
@@
Compared with that for graphical lasso (2.1), KKT condition (2.3) only involves the inverse of the diagonal matrix $\boldsymbol{\Theta}_{D}$, which is trivial to compute. While the second term in (2.3) costs two $p\times p$ matrix-matrix multiplications with $O(p^{3})$ operations, they are much easier to distribute and parallelize than matrix inversion. The sparsity of the optimization variable $\boldsymbol{\Theta}$ imposed by the $\ell_{1}$ penalty further reduces the complexity of the multiplication. Koanantakool et al. (2018) leverage these features to achieve massive scalability in distributed memory HPC systems. This HP-CONCORD implementation can handle 320,000 dimensional data drawn from a random graph structure resembling multi-omic networks.
@@
Despite its computational success, a major drawback of CONCORD is its lack of statistical consistency in estimating $\boldsymbol{\Theta}^{*}$, which is important in downstream tasks in multi-omic analyses as it carries information on the strength of network edges. The objective function in (2.2) defines a loss function $L(\boldsymbol{\Theta},x)=-\log\det\boldsymbol{\Theta}_{D}+\frac{1}{2}\operatorname{\bf tr}(x^{T}\boldsymbol{\Theta}^{2}x)$ for a sample $x\in\mathbb{R}^{p}$. The population risk $R(\boldsymbol{\Theta})=\operatorname{\bf E}L(\boldsymbol{\Theta},X)$ is minimized by a $\boldsymbol{\Theta}$ satisfying the equation $\boldsymbol{\Theta}_{D}^{-1}-\frac{1}{2}(\boldsymbol{\Theta}\boldsymbol{\Sigma}^{*}+\boldsymbol{\Sigma}^{*}\boldsymbol{\Theta})=\mathbf{0}$, where $\boldsymbol{\Sigma}^{*}={\boldsymbol{\Theta}^{*}}^{-1}$ is the covariance matrix of random vector $X$. Unfortunately, the true precision matrix $\boldsymbol{\Theta}^{*}$ does not minimize the risk unless $\boldsymbol{\Theta}_{D}$ equals the identity. This is a very restrictive assumption indeed, and likely does not hold in many applications. Consequently, the currently available result on the consistency of the CONCORD estimator (Theorem 2, Khare, Oh and Rajaratnam, 2015) requires an accurate estimator $\widehat{\boldsymbol{\Theta}_{D}}$ of $\boldsymbol{\Theta}^{*}_{D}$ with a rate $\max_{i=1,\dotsc,p}|\widehat{\boldsymbol{\Theta}_{D}}_{,ii}-\boldsymbol{\Theta}^{*}_{D,ii}|=O_{P}(\sqrt{\log n/n})$. Similarly, SPACE also requires a separate estimator of $\boldsymbol{\Theta}_{D}^{*}$ with the same rate. Such a separate estimator is difficult to find in practice.
@@
There is also an issue with the optimization method employed by HP-CONCORD. This CONCORD-ISTA algorithm (Oh et al., 2014) possesses an $O(1/t)$ rate of convergence, where $t$ is the number of iterations, and the convergence is measured in terms of the objective function value. In addition to the slow, sublinear convergence of the objective function, the variable iterate $\{\boldsymbol{\Theta}^{(t)}\}$ may converge to the minimizer $\hat{\boldsymbol{\Theta}}$ of (2.2) in an arbitrarily slow rate (Bauschke and Combettes, 2011). Since the algorithm has to be terminated within a finite number of iterations, statistical error of $\boldsymbol{\Theta}^{(t)}$ (say $\|\boldsymbol{\Theta}^{(t)}-\boldsymbol{\Theta}^{*}\|$) may remain quite large compared to that of $\hat{\boldsymbol{\Theta}}$ (say $\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{*}\|)$ even after a large number of iterations.
@@ 3.1 ACCORD loss and estimator
For the ACCORD estimator (3.2) to be well-defined, a solution to (3.3) needs to be unique. A matrix $\boldsymbol{X}\in\mathbb{R}^{n\times p}$ $(p\geq n)$ is said to have columns in general position if the affine span of any $n$ points $\{s_{1}X_{i_{1}},\cdots,s_{n}X_{i_{n}}\}$, for arbitrary signs $s_{1},\cdots,s_{n}\in\{-1,1\}$, does not contain any element of {±$\{\pm X_{i}:i\neq i_{1},\cdots,i_{n}\}$. General position occurs almost surely if $\boldsymbol{X}$ is drawn from a continuous distribution (Tibshirani, 2013).
@@ 3.2 ACCORD-FBS algorithm
For the computation of the ACCORD estimator, we use forward-backward splitting (FBS), also known as the proximal gradient method (Combettes and Pesquet, 2007, 2011). Let us split the ACCORD objective function $f$ into $f=g+h$ where
@@
Theory of FBS ensures that with the choice of the step size $\tau_{t}\in(0,2/L)$, where $L=\sigma_{\max}(\boldsymbol{S})$ is the maximum singular value of $\boldsymbol{S}$, the iterate sequence $\{\boldsymbol{\Omega}^{(t)}\}$ converges to the target $\hat{\boldsymbol{\Omega}}$ (Bauschke and Combettes, 2011). The $L$ coincides with the global Lipschitz constant of $\nabla g(\boldsymbol{\Omega})=\boldsymbol{\Omega}\boldsymbol{S}$ in the Euclidean (Frobenius) norm. Backtracking line search adaptively finds a maximal step size $\tau_{t}$ for each iteration $t$ in such a way the convergence is guaranteed (Beck and Teboulle, 2009). Algorithm 1, named ACCORD-FBS, describes both fixed step size and backtracking FBS algorithms for ACCORD. The step size satisfies $\tau_{t}\geq\min\{\tau_{0},\beta/L\}$ for $\beta\in(0,1)$ and the descent property $f(\boldsymbol{\Omega}^{(t+1)})\leq f(\boldsymbol{\Omega}^{(t)})$ is guaranteed by the majorization-maximization (MM) principle (Lange et al., 2022).
@@
Splitting (3.4) is crucial for establishing linear convergence in that the global Lipschitz constant $L<\infty$ of $\nabla g$ exists, and $01/L>$ 0 serves as the lower bound of the step size that guarantees the descent property. Following CONCORD-ISTA (Oh et al., 2014), we may also split $f$ into $\tilde{g}(\boldsymbol{\Omega})=-\log\det\boldsymbol{\Omega}_{D}+({1}/{2})\operatorname{\bf tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{S})$ and $\tilde{h}(\boldsymbol{\Omega})=\lambda\|\boldsymbol{\Omega}\|_{1}$ and call it ACCORD-ISTA. The problem with this splitting is that $\nabla\tilde{g}$ is not globally Lipschitz. The consequence is that the resulting step sizes satisfying the descent condition can be arbitrarily small, contributing to slow convergence of the iterates; see §2 and §5.

### v2

@@ Abstract
Keywords: Ultrahigh-dimensional molecular data, Multi-modal data, High-performance statistical computing, Graphical model selection, Pseudolikelihood, Communication-avoiding linear algebra
@@ 1 Introduction
Omics technologies are essential discovery tools for the analysis of complex biological samples such as tissues and biofluids. Individual omics modalities can generate high-dimensional data (large $p$) for characterizing complex biological samples in a high-throughput manner (large $n\ll p$), and recently for deconvoluting molecular profiles of isolated single cells or spatially connected micrometer-sized regions over tissue sections [37]. Given the unprecedented opportunity to assemble multi-omic data as fine-resolution descriptors of gene expression regulation, cellular metabolism and signal transduction in a biological system, clinical and basic research studies routinely employ a combination of omics platforms to characterize joint variation in the genome, epigenome, transcriptome, proteome and metabolome [12, 16].
@@
In omics-scale analysis, biological network is typically represented as an adjacency matrix of data features based on marginal pairwise correlations. In the popular weighted gene co-expression network analysis [WGCNA, 41], for example, Pearson correlation-based network modules are captured first and the initial network is refined by topological analysis. Despite the success of WGCNA and other extensions in identifying gene communities from omics data, the adjacency of two molecular features as determined by marginal correlations can be a byproduct of shared regulatory factors rather than direct biological interaction. Hence it is natural to estimate the co-expression network underlying a gene expression dataset with a graph representing the conditional dependencies, e.g., partial correlations. The same logic extends to the inference of other types of biological networks from multi-omic data such as metabolic pathways and signaling cascades.
@@
It is well known that the precision matrix, or inverse covariance matrix, of a multivariate distribution encodes partial correlations. The most popular high-dimensional precision matrix estimation approach is the graphical lasso [40, 11], which maximizes a multivariate Gaussian likelihood with an $\ell_{1}$-penalty. However, at the scale of modern multi-omics datasets, it is computationally infeasible to estimate desired networks using this strategy mainly due to the computational bottleneck, which is not easily resolved by using a more powerful computer. As we demonstrate later, our application data features close to 300,000 variables. On two high-performance computing (HPC) environments, the most scalable implementation of graphical lasso, BigQUIC [14], could complete the computation only for the trivial cases in which the computed precision matrix estimates were diagonal. Similarly, a fast implementation [30] of CLIME [4], another popular methodology, was also not applicable on a HPC machine with 192GB of memory when the dataset dimension exceeded 30,000. In the next section, we discuss in detail that the inherent nature of the optimization problems from these popular methods makes it difficult to scale up to modern omics-scale problems.
@@
Feature screening procedures, which reduces the number of variables before the network estimation step, have been proposed as a remedy to address the lack of scalability of methods mentioned above [24, 42]. Although feature screening could be suitable when a moderate dimensionality reduction is necessary, an order of magnitude reduction (300,000 or more to 30,000 or fewer features) precludes functionally important molecules from being considered for network estimation for computational tractability rather than a biologically motivated rationale. It is also difficult to discern essential variables and dispensable ones a priori solely based on numerical criteria chosen by an analyst, hence there is a clear merit in estimating the model with all possible input variables included from the start.
@@
Another category of graphical model structure learning approach is based on optimizing an $\ell_{1}$-penalized pseudolikelihood-based objective function [32, 18]. Although the structure of pseudolikelihood-based objectives can be advantageous from an optimization perspective [20], existing estimators do not possess the statistical properties desirable for multi-omics analyses, as we also discuss in the next section.
@@ 2 Challenges
where $\partial\|\boldsymbol{\Theta}\|_{1}$ denotes the subdifferential of the convex function $\boldsymbol{X}\mapsto\|\boldsymbol{X}\|_{1}$ at $\boldsymbol{\Theta}$. The vast literature on the algorithms for graphical lasso, e.g., [8, 11, 23, 13, 15], essentially reduces to how to solve the KKT equation (2.1) iteratively.
@@
Consequently, most existing algorithms resort to employing an (inner) iterative method to compute $\boldsymbol{\Theta}^{-1}$. For instance, Friedman et al. [11] invoke a $p$-dimensional lasso regression solver $p$ times every (outer) iteration to compute one column of $\boldsymbol{\Theta}^{-1}$ at a time, which becomes prohibitively expensive in omics-scale problems. BigQUIC [14] updates $\boldsymbol{\Theta}$ by block coordinate descent and then computes $\boldsymbol{\Theta}^{-1}$ column by column, by storing only a small fraction of the $p$ columns of $\boldsymbol{\Theta}^{-1}$ in a cache and recomputing the missing columns on demand; the recomputation solves a fraction of the $p$ linear equations defining the inverse ($\boldsymbol{\Theta}\boldsymbol{X}=\boldsymbol{I}_{p}$) by using conjugate gradient. Since this strategy is only successful if the cache miss rate is low, BigQUIC partitions $\boldsymbol{\Theta}$ into a block matrix and permutes blocks to minimize the number of the off-block-diagonal elements. Therefore, the scalability of BigQUIC is limited by the maximum degree of the underlying graph. For more complex graph structures of interest, e.g., those that arise in multi-omics studies, computable dimensions are practically less than $100,000$; see §5.
CLIME minimizes $\|\boldsymbol{\Theta}\|_{1}$ subject to the constraint $\|\boldsymbol{S}\boldsymbol{\Theta}-\boldsymbol{I}_{p}\|_{\infty}\leq\lambda$ over $\mathbb{R}^{p\times p}$, where $\|\cdot\|_{\infty}$ is the vector $\ell_{\infty}$ norm. The optimization problem of CLIME can be decomposed into $p$ independent, $p$-dimensional linear programming (LP) problems, each solving a column of $\boldsymbol{\Theta}$. While LP is the most extensively studied convex optimization problem, commercial LP solvers like Gurobi or Cplex, which utilize interior-point or simplex methods, struggle with scalability when handling hundreds of thousands of variables [25]. Exacerbating the situation, for omics-scale CLIME we are required to solve hundreds of thousands of large-scale LP problems. This remains an almost unattainable task despite the independence. As a result, fastclime [30, 31], which could not handle more than 30,000 variables in our problem instance, remains one of the most scalable implementations publicly available.
As the most scalable pseudolikelihood-based method up to date, CONCORD [18] solves the $\ell_{1}$-penalized minimization problem
@@
where $\boldsymbol{\Theta}_{D}$ denotes the diagonal and $\boldsymbol{\Theta}_{X}$ denotes the off-diagonal parts of $\boldsymbol{\Theta}$. It is a convex amendment of a non-convex pseudolikelihood-based method SPACE [32], which in turn extends the node-wise regression approach by Meinshausen and Bühlmann [27] to estimate $\boldsymbol{\Theta}^{*}$. [^2] CLIME can be understood as an extension of the latter in a different direction, with the aforementioned scalability bottleneck. The associated KKT condition is
@@
Compared with that for graphical lasso (2.1), KKT condition (2.3) only involves the inverse of the diagonal matrix $\boldsymbol{\Theta}_{D}$, which is trivial to compute. While the second term in (2.3) costs two $p\times p$ matrix-matrix multiplications with $O(p^{3})$ operations, they are much easier to distribute and parallelize than matrix inversion. The sparsity of the optimization variable $\boldsymbol{\Theta}$ imposed by the $\ell_{1}$ penalty further reduces the complexity of the multiplication. Koanantakool et al. [20] leverage these features to achieve massive scalability in distributed memory HPC systems. This HP-CONCORD implementation can handle 320,000 dimensional data drawn from a random graph structure resembling multi-omic networks.
@@
Despite its computational success, a major drawback of CONCORD is its lack of statistical consistency in estimating $\boldsymbol{\Theta}^{*}$, which is important in downstream tasks in multi-omic analyses as it carries information on the strength of network edges. The objective function in (2.2) defines a loss function $L(\boldsymbol{\Theta},x)=-\log\det\boldsymbol{\Theta}_{D}+\frac{1}{2}\operatorname{\bf tr}(x^{T}\boldsymbol{\Theta}^{2}x)$ for a sample $x\in\mathbb{R}^{p}$. The population risk $R(\boldsymbol{\Theta})=\operatorname{\bf E}L(\boldsymbol{\Theta},X)$ is minimized by a $\boldsymbol{\Theta}$ satisfying the equation $\boldsymbol{\Theta}_{D}^{-1}-\frac{1}{2}(\boldsymbol{\Theta}\boldsymbol{\Sigma}^{*}+\boldsymbol{\Sigma}^{*}\boldsymbol{\Theta})=\mathbf{0}$, where $\boldsymbol{\Sigma}^{*}={\boldsymbol{\Theta}^{*}}^{-1}$ is the covariance matrix of random vector $X$. Unfortunately, the true precision matrix $\boldsymbol{\Theta}^{*}$ does not minimize the risk unless $\boldsymbol{\Theta}_{D}$ equals the identity. This is a very restrictive assumption indeed, and likely does not hold in many applications. Consequently, the currently available result on the consistency of the CONCORD estimator [Theorem 2, 18] requires an accurate estimator $\widehat{\boldsymbol{\Theta}_{D}}$ of $\boldsymbol{\Theta}^{*}_{D}$ with a rate $\max_{i=1,\dotsc,p}|\widehat{\boldsymbol{\Theta}_{D}}_{,ii}-\boldsymbol{\Theta}^{*}_{D,ii}|=O_{P}(\sqrt{\log n/n})$. Similarly, SPACE also requires a separate estimator of $\boldsymbol{\Theta}_{D}^{*}$ with the same rate. Such a separate estimator is difficult to find in practice.
@@
There is also an issue with the optimization method employed by HP-CONCORD. This CONCORD-ISTA algorithm [29] possesses an $O(1/t)$ rate of convergence, where $t$ is the number of iterations, and the convergence is measured in terms of the objective function value. In addition to the slow, sublinear convergence of the objective function, the variable iterate $\{\boldsymbol{\Theta}^{(t)}\}$ may converge to the minimizer $\hat{\boldsymbol{\Theta}}$ of (2.2) in an arbitrarily slow rate [2]. Since the algorithm has to be terminated within a finite number of iterations, statistical error of $\boldsymbol{\Theta}^{(t)}$ (say $\|\boldsymbol{\Theta}^{(t)}-\boldsymbol{\Theta}^{*}\|$) may remain quite large compared to that of $\hat{\boldsymbol{\Theta}}$ (say $\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{*}\|)$ even after a large number of iterations.
@@ 3.1 ACCORD loss and estimator
For the ACCORD estimator (3.2) to be well-defined, a solution to (3.3) needs to be unique. A matrix $\boldsymbol{X}\in\mathbb{R}^{n\times p}$ $(p\geq n)$ is said to have columns in general position if the affine span of any $n$ points $\{s_{1}X_{i_{1}},\cdots,s_{n}X_{i_{n}}\}$, for arbitrary signs $s_{1},\cdots,s_{n}\in\{-1,1\}$, does not contain any element of {±$\{\pm X_{i}:i\neq i_{1},\cdots,i_{n}\}$. General position occurs almost surely if $\boldsymbol{X}$ is drawn from a continuous distribution [35].
@@ 3.2 ACCORD-FBS algorithm
For the computation of the ACCORD estimator, we use forward-backward splitting (FBS), also known as the proximal gradient method [6, 7]. Let us split the ACCORD objective function $f$ into $f=g+h$ where
@@
Theory of FBS ensures that with the choice of the step size $\tau_{t}\in(0,2/L)$, where $L=\sigma_{\max}(\boldsymbol{S})$ is the maximum singular value of $\boldsymbol{S}$, the iterate sequence $\{\boldsymbol{\Omega}^{(t)}\}$ converges to the target $\hat{\boldsymbol{\Omega}}$ [2]. The $L$ coincides with the global Lipschitz constant of $\nabla g(\boldsymbol{\Omega})=\boldsymbol{\Omega}\boldsymbol{S}$ in the Euclidean (Frobenius) norm. Backtracking line search adaptively finds a maximal step size $\tau_{t}$ for each iteration $t$ in such a way the convergence is guaranteed [3]. Algorithm 1, named ACCORD-FBS, describes both fixed step size and backtracking FBS algorithms for ACCORD. The step size satisfies $\tau_{t}\geq\min\{\tau_{0},\beta/L\}$ for $\beta\in(0,1)$ and the descent property $f(\boldsymbol{\Omega}^{(t+1)})\leq f(\boldsymbol{\Omega}^{(t)})$ is guaranteed by the majorization-maximization (MM) principle [22].
@@
Splitting (3.4) is crucial for establishing linear convergence in that the global Lipschitz constant $L<\infty$ of $\nabla g$ exists, and $01/L>$ 0 serves as the lower bound of the step size that guarantees the descent property. Following CONCORD-ISTA [29], we may also split $f$ into $\tilde{g}(\boldsymbol{\Omega})=-\log\det\boldsymbol{\Omega}_{D}+({1}/{2})\operatorname{\bf tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{S})$ and $\tilde{h}(\boldsymbol{\Omega})=\lambda\|\boldsymbol{\Omega}\|_{1}$ and call it ACCORD-ISTA. The problem with this splitting is that $\nabla\tilde{g}$ is not globally Lipschitz. The consequence is that the resulting step sizes satisfying the descent condition can be arbitrarily small, contributing to slow convergence of the iterates; see §2 and §5.